One side of triangle is
2
2
times the shortest side. The third side is
24
24
feet more than the shortest side. The perimeter is
84
84
. Find all three sides.

Answer by ikleyn(52879)

.

    x + 2x + (x+24) = 84.


    4x + 24 = 84

    4x      = 84 - 24

    4x      = 60

     x       = 60/4 = 15.


The sides should be  15, 30 and 39 units.    ANSWER


CHECK.  The found side lengths satisfy the triangle inequalities --- hence, such a triangle does exist.


ANSWER.  The triangle side lengths, obtained from the condition, are 15, 30 and 39 units.

Solved, explained and checked.